# Topological interpretation of extremal and Davies-type phase transitions of black holes

### v1

@@ 1 Introduction
Over the years the phase transition of black holes has been studied quite extensively. Earlier, the thermogeometrical interpretation of phase transitions was provided [16, 28, 29, 30, 31, 32] (also see the unified approaches [33, 34, 35]), where the critical points were identified as ones that are endowed with the diverging Ricci-scalar in the thermodynamic phase space. This (thermogeometrical) interpretation has been found to hold good for all types of phase transitions described above. Recently, the topological interpretation has been provided by researchers, which helps to distinguish the properties of the black holes, without resorting to the dynamical equation of the spacetime or the structural details. This formalism was earlier used in the coordinate space of black holes [36, 37], to identify the location of the light rings using null geodesics. It was shown that the location of the light rings correspond to the zero point of a special vector, constructed in the coordinate space of the black hole spacetime. Subsequently, for a black hole in asymptotically flat spacetime with a topologically spherical Killing horizon, it was found that the black holes have at least one light ring outside the horizon for each sense of rotation [37], which was later later generalized to other cases [38, 39, 40]. Later, this topological argument was used in different aspects of black hole thermodynamics [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62]. In the black hole phase transition, it was shown that the critical point can have a topological interpretation in terms of non-zero topological charge [42, 43, 44, 45, 46, 47, 48, 49]. So far, this interpretation has been found to hold for the van der Waals type phase transition [42, 43, 44, 45, 46] and the Hawking-Page phase transition [47, 48, 49]. It was first obtained for the van der Waals type phase transition [42], where it was shown that a topological interpretation can be provided using Duan’s topological current $\phi$-mapping theory [63, 64]. The same analysis has been extended later to the Hawking-Page transition. However, the topological interpretation is still missing for the other two types of phase transitions in black holes, namely the Davies type phase transition and the extremal phase transition. Therefore, it becomes an obvious question whether the same interpretation exists for these other two types of phase transitions. In this paper, we will explore this issue, i.e. we obtain a topological interpretation for the Davies phase transition and for the extremal phase transition. The entire analysis is general as it is valid for any spacetime which is endowed with these types of phase transitions.
@@ 5 Conclusions:
Inspired by [42], where the topological interpretation has been provided for van der Waals-type phase transition, we have adopted Duan’s $\phi$-mapping topological current theory to provide the topological interpretation for Davies-type and extremal phase transitions. Thereby, we show that the topological interpretation can indeed be provided for these two types of phase transitions by making judicious choices of thermodynamic potential $\boldsymbol{\Phi}$ in every case. For the Davies-type phase transition, which is characterized by the divergence of heat capacity, the inverse Hawking temperature ($T^{-1}$) plays the role of the thermodynamic potential to obtain the topological interpretation. On the other hand, for the extremal phase transition, we found that the mass of the black hole plays the role of thermodynamic potential. In both cases, we have chosen the thermodynamic potential in such a way that the critical point lies on the zero-point of the vector field $\phi^{a}$, which is obtained from the thermodynamic potential $\boldsymbol{\Phi}$. Despite the entire description being provided from a general standpoint, we also have provided specific examples to support our claim and have shown that the topological argument holds good for the black hole phase transition. In all cases, we find that the critical point can be endowed with non-zero topological charges using the $\phi$-field formalism. More importantly, we show that the recent development of topological interpretation of black hole phase transitions, which has been so far explored for van der Waals and Hawking-Page phase transition, can be extended for other criticalities as well. Thus, the topological interpretation can act as a standard formalism to describe the phase transition in black holes.
Finally, as we have mentioned earlier, the topological argument is provided in order to describe the properties of the black holes without involving the field equation and detail structure. Hence, the formalism mentioned here can be applicable beyond Einstein’s GR and to any black holes that show these types of phase transitions.

### v2

@@ 1 Introduction
Over the years the phase transition of black holes has been studied quite extensively. Earlier, the thermogeometrical interpretation of phase transitions was provided [16, 28, 29, 30, 31, 32] (also see the unified approaches [33, 34, 35]), where the critical points were identified as ones that are endowed with the diverging Ricci-scalar in the thermodynamic phase space. This (thermogeometrical) interpretation has been found to hold good for all types of phase transitions described above. Recently, the topological interpretation has been provided by researchers, which helps to distinguish the properties of the black holes, without resorting to the dynamical equation of the spacetime or the structural details. This formalism was earlier used in the coordinate space of black holes [36, 37], to identify the location of the light rings using null geodesics. It was shown that the location of the light rings correspond to the zero point of a special vector, constructed in the coordinate space of the black hole spacetime. Subsequently, for a black hole in asymptotically flat spacetime with a topologically spherical Killing horizon, it was found that the black holes have at least one light ring outside the horizon for each sense of rotation [37], which was later later generalized to other cases [38, 39, 40]. Later, this topological argument was used in different aspects of black hole thermodynamics [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68] (also see the recent papers [69, 70, 71, 72]). In the black hole phase transition, it was shown that the critical point can have a topological interpretation in terms of non-zero topological charge [42, 43, 44, 45, 46, 47, 48, 50]. So far, this interpretation has been found to hold for the van der Waals type phase transition [42, 43, 44, 45, 46] and the Hawking-Page phase transition [47, 48, 50]. It was first obtained for the van der Waals type phase transition [42], where it was shown that a topological interpretation can be provided using Duan’s topological current $\phi$-mapping theory [73, 74]. The same analysis has been extended later to the Hawking-Page transition. However, the topological interpretation is still missing for the other two types of phase transitions in black holes, namely the Davies type phase transition and the extremal phase transition. Therefore, it becomes an obvious question whether the same interpretation exists for these other two types of phase transitions. In this paper, we will explore this issue, i.e. we obtain a topological interpretation for the Davies phase transition and for the extremal phase transition. The entire analysis is general as it is valid for any spacetime which is endowed with these types of phase transitions.
@@ 5 Conclusions:
Recently, it was found that the thermodynamic properties of black holes can be deduced when one uses a phenomenological theory of topology in the thermodynamic phase space. At first [41], it was shown that all the black hole solutions can be identified in three groups based on the topological charges defined in the thermodynamic parameter space i.e. $Q=-1,0,$ and $+1$. In addition, it was shown that the black holes with the same topological nature exhibit similar thermodynamic properties. Later, this topological argument has been extended to explore the phase transitions in black holes. Firstly, it was shown in [42] that the critical point in the Van der Waals type phase transition of the black holes can have a topological interpretation using Duan’s $\phi$-field formalism in the thermodynamic parameter space. In that case, the critical point can be attributed to a non-zero topological charge. Later, this argument has been extended to the Hawking-Page phase transition [47]. However, the topological interpretation is still missing in literature for the other two types of phase transitions i.e. the Davies-type phase transition and the extremal phase transition. Thus, inspired by [42], we have adopted Duan’s $\phi$-mapping topological current theory to provide the topological interpretation for Davies-type and extremal phase transitions. Thereby, we show that the topological interpretation can indeed be provided for these two types of phase transitions by making judicious choices of the thermodynamic potential $\boldsymbol{\Phi}$ in every case. For the Davies-type phase transition, which is characterized by the divergence of heat capacity, the inverse Hawking temperature ($T^{-1}$) plays the role of the thermodynamic potential to obtain the topological interpretation. On the other hand, for the extremal phase transition, we found that the mass of the black hole plays the role of the thermodynamic potential. In both cases, we have chosen the thermodynamic potential in such a way that the critical point lies on the zero-point of the vector field $\phi^{a}$, which is obtained from the thermodynamic potential $\boldsymbol{\Phi}$. Despite the entire description being provided from a general standpoint, we also have provided specific examples to support our claim and have shown that the topological argument holds good for the black hole phase transition. In all cases, we find that the critical point can be endowed with non-zero topological charges using the $\phi$-field formalism. More importantly, we show that the recent development of topological interpretation of black hole phase transitions, which has been so far explored for van der Waals and Hawking-Page phase transition, can be extended for other criticalities as well. Thus, the topological interpretation can act as a standard formalism to describe the phase transition in black holes.
Finally, as we have mentioned earlier, the topological argument is provided in order to describe the properties of the black holes without involving the field equation and detail structure. Hence, the formalism mentioned here can be applicable beyond Einstein’s GR and to any black holes that show these types of phase transitions. Let us explain it in a bit more detail. The thermodynamic properties of black holes were initially established in general relativity [1, 2, 3]. However, later it was found that the thermodynamic properties are valid well beyond general relativity [81, 82, 83, 84, 85, 87, 86, 88, 89, 90] (for a recent review on this matter, see [91]). Phase transition, which is a key aspect of black hole thermodynamics, has been found beyond GR as well. For instance, Van der Waals-type phase transition has been observed in higher curvature gravity and scalar-tensor gravity [92, 93, 94, 95]. Similarly, Davies-type phase transition has been observed beyond general relativity such as in Horava-Lifshitz gravity [96], Gauss-Bonnet gravity [97], Lovelock gravity [98], $f(R)$ gravity [99], and so on. Furthermore, one can also find extremal black hole solutions beyond GR [100] and one can expect to find extremal phase transition beyond general relativity as well. We emphasize that our analysis will continue to be valid for all those cases as our topological argument has been formulated in the thermodynamic parameter space and does not involve any gravitational equation.
Since our topological argument is formulated in the thermodynamic parameter space of a black hole (BH), we admit that the observational connection with our analysis is not direct and we need to investigate much deeper to comment on the potential connection with black hole shadow, accretion disc etc. However, our analysis distinguishes certain black holes from others, which can have significance from the observational viewpoint. For example, it distinguishes extremal black holes (non-zero topological charge) from the non-extremal ones. Therefore, the extremal black holes can be identified from our analysis. Furthermore, we discussed that for Davies-type phase transition, heat capacity diverges for certain black holes at a specific horizon radius. A closer investigation (see [101] for example) shows that the larger BHs show positive heat capacity and the smaller BHs show negative heat capacity and at the critical limit, heat capacity diverges. Therefore, smaller BHs are relatively unstable. From our topological argument, one can obtain the stable limit (non-zero topological charge) of different black holes. To obtain further connection with other observable features, we need to investigate further, which we keep as a potential future work.